OpenStudy (anonymous):

How do I change (1,4) and slope=-2 into a slope intercept form

Mehek (mehek14):

slope intercept form y = mx + b

OpenStudy (qwertty123):

Yup^

OpenStudy (anonymous):

Yes

Mehek (mehek14):

m = slope b = y-intercept

Mehek (mehek14):

you already have the slope -2 put that instead of m

OpenStudy (anonymous):

Ok then what

OpenStudy (qwertty123):

@Mehek14 You go!

Mehek (mehek14):

y = -2x + b now plug in the ordered pari x = 1 y = 4 4 = -2*1 + b solve for b

OpenStudy (anonymous):

So 2

Mehek (mehek14):

4 = -2*1 + b 4 = -2 + b 6 = b it's 6 not 2

OpenStudy (anonymous):

Oh I messed up

Mehek (mehek14):

it's fine so b = 6 y = -2x + 6

OpenStudy (anonymous):

So what about (-4,6) and slope=-2/3

Mehek (mehek14):

well we got slope -2/3 y = -2/3x + b 6 = -2/3 * 4 + b solve for b

OpenStudy (anonymous):

So what about (-4,6) and slope=-2/3

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 well we got slope -2/3 y = -2/3x + b 6 = -2/3 * 4 + b solve for b \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (anonymous):

So it's 5

Mehek (mehek14):

no b = 8.666

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

So its not a whole number

Mehek (mehek14):

nope

OpenStudy (anonymous):

So I have (-5,-11) , (-2,-1 )

Mehek (mehek14):

what about those ordered pairs?

OpenStudy (anonymous):

How do i make then into a slope intercept

Mehek (mehek14):

first find the slope \(\dfrac{-11 - (-1)}{-5-(-2)}\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

-12/-7

OpenStudy (anonymous):

No its -10/-3

Mehek (mehek14):

ye -10/-3

Mehek (mehek14):

-10/-3 = 10/3 so you have y = 10/3x + b

OpenStudy (anonymous):

What's the b

Mehek (mehek14):

y-intercept

OpenStudy (anonymous):

What's that

Mehek (mehek14):

the place where the line goes through the y-axis

OpenStudy (anonymous):

I have y=-10/-3+ blank

OpenStudy (anonymous):

For my answer

Mehek (mehek14):

so we plug in one of the ordered pairs into the equation

OpenStudy (anonymous):

So I can put 1 there

Mehek (mehek14):

use (-2,-1) -2 = 10/3*-1 + b solve for b

OpenStudy (anonymous):

Is it a decimal

Mehek (mehek14):

ye

OpenStudy (anonymous):

Is it -5.3

OpenStudy (anonymous):

You can do point slope form. So you do: y-4=-2(x-1) y-4=-2x + 2 y = -2x + 6